OpenStudy (anonymous):

Suppose the number of babies born in Hong Kong in 1994 is 70000 and in subsequent years, the number of babies born each year increases by 2% of that of the previous year. (a) Find the number of babies born in HK (i) in the first year after 1994. [71400] (ii) in the nth year after 1994. [70000*1.02^(n)] (b) In which year will the number of babies born in HK first exceed 90000? [n=12] >>>>(c) Find the total number of babies born in HK from 1997 to 2046 inclusive.<<<< >>>>(d) It is known that from 1901 to 2099, a year is a leap year if its number is divisible by 4. (i) Find the number of leap years between 1997 and 2046. (ii) Find the total number of babies born in Hong Kong in the leap years between 1997 and 2046.<<<<

OpenStudy (anonymous):

@Directrix @Hero

OpenStudy (anonymous):

\(100\%+2\%=1.02\)

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

\(1.2^n\)?

OpenStudy (anonymous):

yes, it's a geometric sequence

OpenStudy (anonymous):

Where is \(1.2\) from? I only see \(1.02\)

OpenStudy (anonymous):

ohh, i type it wrongly.

OpenStudy (anonymous):

my problem is in part c

OpenStudy (anonymous):

Okay, first find \(n\)

OpenStudy (anonymous):

\[ \sum_{k=0}^{n}ar^k=a\frac{1-r^{n+1}}{1-r} \]

OpenStudy (anonymous):

ohh, wait

OpenStudy (anonymous):

n=12

OpenStudy (anonymous):

\[ n=2046-1997 \]

OpenStudy (anonymous):

The only problem is that \(k\) starts at \(4\)

OpenStudy (anonymous):

ohh, you are talking about part c @@ n=49

OpenStudy (anonymous):

what's k?

OpenStudy (anonymous):

I mean \(3\)

OpenStudy (anonymous):

i am thinking there are 49years or 50 years..

OpenStudy (anonymous):

Okay let's actually do \(n_1= 1997-1994\) and \(n_2=2046-1994\)

OpenStudy (anonymous):

\[ \sum_{k=0}^{n_2}ar^k -\sum_{k=0}^{n_1}ar^k \]

OpenStudy (anonymous):

Basically it is telling you how to add up geometric series.

OpenStudy (anonymous):

okk...

OpenStudy (anonymous):

So basically add up all children from 1994 to 2046 and then subtract all children from 1994 to 1996

OpenStudy (anonymous):

oh, i see

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio \[ \sum_{k=0}^{n}ar^k=a\frac{1-r^{n+1}}{1-r} \] \(\color{blue}{\text{End of Quote}}\) This is the formula for adding up form year 0 to year n

OpenStudy (anonymous):

year 0 is 1994

OpenStudy (anonymous):

70000*(1+2%)^2=72828 70000*(1+2%)^52=196022.973 196022.973-72828=123194.973 ?!

OpenStudy (anonymous):

No, look at the formula. You only calculating for a particular year.

OpenStudy (anonymous):

what do you mean..

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

You need to calculate the sum of each year.

OpenStudy (anonymous):

ohh my god, so complicated

OpenStudy (anonymous):

You could at least try the formula I gave.

OpenStudy (anonymous):

I mean... we know \(a=70000\), \(r=1.02\)

OpenStudy (anonymous):

S(52)=6301148.649

OpenStudy (anonymous):

Going from 1994 to 1996 we have \(n=2\) Going from 1994 to 2046 we have \(n=52\)

hartnn (hartnn):

find the sum from 1994 to 2046 and subtract that with the sum from 1994 to 1997 you will get sum for 1997 to 2046

OpenStudy (anonymous):

\[ 70000\frac{1-(1.02)^{52+1}}{1-1.02}- 70000\frac{1-(1.02)^{2+1}}{1-1.02} \]

hartnn (hartnn):

i see thats already been proposed

OpenStudy (anonymous):

why +1?

OpenStudy (anonymous):

Because the formula has \(n+1\) in it.

OpenStudy (anonymous):

not n-1?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio \[ \sum_{k=0}^{n}ar^k=a\frac{1-r^{n+1}}{1-r} \] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

how about part d?

OpenStudy (anonymous):

The first problem is division

OpenStudy (anonymous):

eg 1904, 2096 are leap years?

OpenStudy (anonymous):

The second problem requires that you create a new formula which is just number of children in a leap year.

OpenStudy (anonymous):

So the new \(r\) is going to be \(1.02^4\)

OpenStudy (anonymous):

i got 27 lol

OpenStudy (anonymous):

@ganeshie8

OpenStudy (phi):

(i) Find the number of leap years between 1997 and 2046. if we start and end on a leap year, for example 2000 to 2008 (where we know we have 3 leap years) we do the following : 2008 - 2000 = 8 years divide by 4, to get 2 add 1 to get 3 leap years between 2000 and 2008 inclusive. you can test other years, and see the formula is (last_year - first_year) /4 + 1 as long as both the last_year and first_year are leap years. in your case 1997 and 2046. the first leap year in this interval is 2000 ( you can use the rule: if the last 2 digits are evenly divisible by 4, the whole number is divisible by 4) the last leap year in this range is 2044 (44 is evenly divisible by 4) now use the formula: 2044- 2000 = 44 divide by 4 to get 11 add 1 we find that there are 12 leap years between 1997 and 2046.

OpenStudy (phi):

(ii) Find the total number of babies born in Hong Kong in the leap years between 1997 and 2046. from part (i), we know the range will be leap years from 2000 to 2044 inclusive We can use the result of (a, ii) to find the number born in the year 2000 call this number P2000 the number born in "0th leap yr" is P2000 the number born in 2004 "1st leap yr" P2000 * (1.02)^4 the number born in 2008 "2nd leap yr" P2004* (1.02)^4 or P2000* (1.02^4)^2 if you carry out this pattern you see the number born in the nth leap year is P2000* (1.02^4)^n if we let 1.02^4 = r then the sum of the those born in leap years from number 0 (yr 2000) to number 11 (yr 2044) is \[\sum_{n=0}^{11} r^n= \frac{1-r^{12}}{1-r}\]

OpenStudy (anonymous):

how come i get 6255546.587?? @phi @hartnn

OpenStudy (anonymous):

@Hero

OpenStudy (phi):

what did you get for (1- r^12)/(1-r) ?

OpenStudy (phi):

I get r= 1.02^4 = 1.0824 (1-r^12)/(1-r) = 19.2530 the population in the year 2000 is 70000*1.02^6 = 78831.37 times 19.253 gives 1,517,744.188

OpenStudy (anonymous):

i got it now. Thanks @phi